OpenStudy (anonymous):

Help! I'm stuck Solve the system using the elimination method. 2x + 3y = 5 8y = x + 26

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@austinL @satellite73

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

I found the x point first. It was -3.8.

OpenStudy (anonymous):

but then I tried to plug it back into the 2nd equation and it didn't work out

OpenStudy (anonymous):

I'll copy paste my work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

First you have to line up the x’s and y’s. 2x+3=5 -x+8y=26 You have to make either the x’s or the y’s cancel out. Looking at our problem, it would be easier to do the x’s. We could multiply 2 to the 2nd equation so that it becomes -2. Then we can cancel 2x and -2x. 2x + 3y = 5 2(-x+8y= 26) --> -2x-16y=52 2x and -2x cancels out. Simplify the rest. 2x + 3y = 5 -2x-16y=52 ________ -13y= 50 Y= -3.8 Now we plug the y value into any equation to get the x value.

OpenStudy (anonymous):

this is where you got stuck

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

when i tried to plug it back into the 2nd equation to find the x value.

OpenStudy (anonymous):

is everything i did correct?

OpenStudy (anonymous):

ok so if it works for one it should work for the other so lets try a different number

OpenStudy (anonymous):

this is when I tried to plug it back in: -x+8y=26 -(-3.8)+8(-3.8) = 26 3.8 + 30.4 = 26 34.2 = 26

OpenStudy (anonymous):

ok. so solve for x first instead of y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u write it out for me I kinda have to be somewhere soon

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try 3.2

OpenStudy (anonymous):

to plug in?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or 3.5

OpenStudy (anonymous):

ohh, i know my mistake, i plugged in the -3.8 for x AND y in the 2nd equation, when i was supposed to only plug it in for y. let me see if it works.

OpenStudy (anonymous):

i got 56.4 for x. is that too big?

OpenStudy (anonymous):

try anything below 3.8

OpenStudy (anonymous):

why

OpenStudy (anonymous):

and then you will get you answer ooo i got somebody who can help

OpenStudy (anonymous):

I got -3.8 because 50 divided by -13 is -3.8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

look! http://www.algebra.com/algebra/college/linear/solving-linear-system-by-elimination.solver where it got worked out

OpenStudy (anonymous):

oh darn it, another person can't see it

OpenStudy (anonymous):

i can see it wht grade are you in+

OpenStudy (anonymous):

look at the screenshot

OpenStudy (anonymous):

2nd page.

OpenStudy (anonymous):

so the answer is (-2, 3)

OpenStudy (anonymous):

yes and your grade

OpenStudy (anonymous):

huh?